Differentiate Sin^2 x with respect to e^ cosx

Grade:

3 Answers

Sajin SS
11 years ago

Dear Aswathi

 we know that dy/dz can be written as

                 dy/dz = (dy/dx) / (dz/dx)

So here, y=sin^2x & z = e^cosx

  so dy/dx = 2 sinx cosx           &          dz/dx = -sinx (e^cosx)

Thus , your answer is

             (dy/dx) / (dz/dx) = - (2cosx / e^cosx)

:)  Hit like if u like the answer :)

Harshal Sureshbhai Davda
11 years ago

I think answer is

=2*(sin x)*e^(-cos x)

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your question.
 
Let, y = sin2x and z = ecosx
dy/dx = 2sinx.cosx  ;    dz/dx = ecosx.( – sinx)
Now, dy/dz = (dy/dx).(dx/dz) = (dy/dx)/(dz/dx)
= 2sinxcosx/( – ecosxsinx)
= – 2cosx.e–cosx
 
Thanks and regads,
Kushagra